# The Racah algebra of rank 2: Properties, symmetries and representation

### v1

@@ Appendix A Appendix — Proofs of double commutators
$[D_{ijk},[D_{jk\ell},P_{jk}]]$ $=$ $[[P_{jk},D_{jk\ell}],D_{ijk}]$
$=$ $[(P_{k\ell}-P_{j\ell})P_{jk}+2D_{jk\ell}-2P_{j}P_{k\ell}+2P_{k}P_{j\ell},D_{ijk}]$
$=$ $(P_{ik}P_{j\ell}+P_{ij}P_{k\ell}-2P_{i\ell}P_{jk})P_{jk}$
$+(P_{k\ell}-P_{j\ell})\left((P_{ik}-P_{ij})P_{jk}+2D_{ijk}-2P_{j}P_{ik}+2P_{k}P_{ij}\right)$
$-2(D_{ji\ell}+D_{ki\ell})P_{jk}-2P_{j}[P_{\ell k},D_{kij}]+2P_{k}[P_{\ell j},D_{jki}]$
$=$ $-2P_{i\ell}P_{jk}^{2}+(P_{k\ell}P_{ik}+P_{j\ell}P_{ij})P_{jk}$
$+2(P_{k\ell}-P_{j\ell})D_{ijk}-2D_{i\ell j}P_{jk}-2D_{i\ell k}P_{jk}$
$-2P_{j}\left(P_{k\ell}P_{ij}-P_{i\ell}P_{jk}\right)+2P_{k}\left(P_{i\ell}P_{jk}-P_{j\ell}P_{ij}\right).$
@@
$+2P_{j}\left(P_{ik}P_{j\ell}-P_{i\ell}P_{jk}\right)-2P_{k}\left(P_{i\ell}P_{jk}-P_{ij}P_{j\ell}\right).$
Notice that the triple of $P^{\prime}s$ in (A.2) and (A.2) will sum together to give $2D_{\ell ki}P_{jk}+2D_{\ell ji}P_{jk}$ which will cancel with subsequent terms.
Putting everything together we arrive at
@@
The first term on the right has not been involved in any of the previous computations as all of the previous computations has decreased the order of terms to at most triples of $P^{\prime}s$, or equivalently a $P$ and $D$. We can remove this term by taking the commutator of a $PD$ relation (2.35) (again with the opposite ordering with a $P$ as in
$[P_{jk},2D_{jk\ell}P_{i}+D_{k\ell i}P_{ij}+D_{\ell ji}P_{ik}+D_{jki}P_{i\ell}]=0,$ (A.7)
@@
$+2D_{k\ell i}D_{kji}+2D_{\ell ji}D_{jki}=0.$
Thus, after a permutation of indices we obtain
@@
$[P_{jk},2D_{jk\ell}P_{i}+D_{k\ell i}P_{ij}+D_{\ell ji}P_{ik}+D_{jki}P_{i\ell}]=In+Out,$
with $In$ and $Out$ a collection of terms involving inner and outer quadratic relations, respectively, that is
$In=-P_{ij}[P_{jk},D_{jk\ell}]+[P_{i\ell}[P_{jk},D_{jki}]]+P_{j\ell}[P_{ik},D_{ikj}]-[P_{ij},D_{ij\ell}]P_{jk},$
$Out$ $=$ $\left([P_{ji},D_{i\ell k}]+[P_{\ell i},D_{ijk}]-[P_{ij},D_{jk\ell}]\right)P_{jk}$
$+\left([P_{jk},D_{ki\ell}]+[P_{ik},D_{jk\ell}]\right)P_{ij}$
$+\left([P_{kj},D_{j\ell i}]+[P_{\ell j},D_{jik}]\right)P_{ik}+P_{ik}[P_{ij},D_{jk\ell}]-2P_{j}[P_{ik},D_{k\ell j}].$
Beginning with the inner relation terms, we can use the lower-order Jacobi identity proven earlier
$[P_{i\ell}[P_{jk},D_{jki}]]=[P_{jk},[P_{i\ell},D_{jki}]],$
to simplify the term in $In$
$[P_{i\ell}[P_{jk},D_{jki}]]$ $=$ $[P_{jk},P_{ij}P_{k\ell}-P_{ik}P_{j\ell}]$ (A.13)
$=$ $P_{jk}P_{ij}P_{k\ell}-P_{jk}P_{j\ell}P_{ik}-(P_{ij}P_{k\ell}-P_{ik}P_{j\ell})P_{jk}$
$=$ $2D_{kji}P_{k\ell}+2P_{ij}D_{jk\ell}-2D_{kj\ell}P_{ik}-2P_{j\ell}D_{jki}$
$=$ $2D_{jik}P_{k\ell}+2D_{\ell jk}P_{ki}+2P_{ij}D_{jk\ell}+2P_{\ell j}D_{jik}.$
The other 3 terms from $In$ give
@@
$=$ $P_{ij}\left(P_{j\ell}P_{jk}-P_{jk}P_{k\ell}-2P_{k}P_{j\ell}+2P_{j}P_{k\ell}\right)$
$+P_{j\ell}\left(P_{ik}P_{jk}-P_{ij}P_{ik}-2P_{i}P_{jk}+2P_{k}P_{ij}\right)$
$+\left(P_{i\ell}P_{ij}-P_{ij}P_{j\ell}-2P_{j}P_{i\ell}+2P_{i}P_{j\ell}\right)P_{jk}$
$=$ $(P_{ik}P_{j\ell}-P_{ij}P_{k\ell})P_{jk}-2P_{ij}D_{jk\ell}$
$+(P_{i\ell}P_{jk}-P_{ik}P_{j\ell})P_{ij}+2P_{i\ell}D_{ijk}-2P_{j\ell}D_{jik}$
$+4P_{k}D_{\ell ji}+2P_{j}(P_{ij}P_{k\ell}-P_{i\ell}P_{jk}).$
@@
Moving on to the $Out$ term. We use Lemma 7 several times to simplify
$Out$ $=$ $[P_{jk},D_{k\ell i}]P_{jk}+[P_{ij},D_{j\ell k}]P_{ij}$
$-[P_{ij},D_{jk\ell}]P_{ik}+P_{ik}[P_{ij},D_{jk\ell}]-2P_{j}[P_{ik},D_{k\ell j}].$
@@
$Out=[P_{jk},D_{k\ell i}]P_{jk}+[P_{ij},D_{j\ell k}]P_{ij}+2D_{ki\ell}P_{jk}+2P_{i\ell}D_{ikj}-2P_{j}[P_{ik},D_{k\ell j}].$ (A.16)
Hence, we have shown $In+Out=0$ and the Jacobi identity holds for these generators without additional constraints.
A.4 Case $\{D_{ij\ell},D_{ijk},D_{jk\ell}\}$
The final case will be for 3 distinct $D^{\prime}s$. Without loss of generality we chose the indices as above and consider the quantity
$[D_{ij\ell},[D_{ijk},D_{jk\ell}]]+[D_{ijk},[D_{jk\ell},D_{ij\ell}]]+[D_{jk\ell},[D_{ij\ell},D_{ijk}]].$ (A.17)
@@ A.4 Case $\{D_{ij\ell},D_{ijk},D_{jk\ell}\}$
Appendix B Appendix — Coefficients of the action of the proposed representation
@@
$n_{M}=N+\sum_{i\in M}c_{i},\qquad M\subseteq\{1,2,3,4\}.$
$C_{1}|t,s\rangle=c_{1}(c_{1}-1)|t,s\rangle\hskip 426.79134pt$
$C_{2}|t,s\rangle=c_{2}(c_{2}-1)|t,s\rangle\hskip 426.79134pt$
$C_{3}|t,s\rangle=c_{3}(c_{3}-1)|t,s\rangle\hskip 426.79134pt$
$C_{4}|t,s\rangle=c_{4}(c_{4}-1)|t,s\rangle\hskip 426.79134pt$
$C_{1234}|t,s\rangle=n_{1234}(n_{1234}-1)|t,s\rangle\hskip 426.79134pt$
[…]
$\varphi_{t,s}=(s-t)(N+1-t)(N+2c_{2}-t)(2n_{123}-t-s-1)$
$\theta_{t,s}=(n_{12}-t)(n_{12}-t-1)$
$C_{23}|t,s\rangle$ $=$ $\theta^{*}_{t,s}|t,s\rangle+|t+1,s\rangle\hskip 341.43306pt$
$\theta^{*}_{t,s}=(n_{23}-t)(n_{23}-t-1)$
$C_{123}|t,s\rangle$ $=$ $\nu_{t,s}|t,s\rangle\hskip 341.43306pt$
$\nu_{t,s}=(n_{123}-s)(n_{123}-s-1)$
$C_{34}|t,s\rangle$ $=$ $\hat{\phi}^{*}_{t,s}|t-1,s+1\rangle+\hat{\vartheta}^{*}_{t,s}|t,s+1\rangle\hskip 341.43306pt$
$+\phi^{*}_{t,s}|t-1,s\rangle+\vartheta^{*}_{t,s}|t,s\rangle$
$+\check{\phi}^{*}_{t,s}|t-1,s-1\rangle+\check{\vartheta}^{*}_{t,s}|t,s-1\rangle$
$\hat{\phi}^{*}_{t,s}=-(s-t)(s-t+1)(N+1-t)(N+2c_{2}-t)$
$\phi^{*}_{t,s}=\varphi_{t,s}(1-\psi_{t,s})$
$\hskip 19.91692pt=\left(\dfrac{1}{2}-\dfrac{n_{123}(n_{123}+2c_{4}-1)}{2(n_{123}-s)(n_{123}-s-1)}\right)$
$\hskip 36.98857pt\cdot(s-t)(N+1-t)(N+2c_{2}-t)(2n_{123}-t-s-1)$
@@
$\hskip 19.91692pt=-s\dfrac{(s+2c_{4}-1)(2n_{123}-s)(2n_{1234}-s-1)(2n_{123}-t-s-1)}{4(2n_{123}-2s+1)(2n_{123}-2s-1)(n_{123}-s)^{2}}$
$\hskip 36.98857pt\cdot(N+1-t)(2n_{123}-t-s)(N+2c_{2}-t)$
$\hat{\vartheta}^{*}_{t,s}=-(s-t)(s-2c_{3}-t+1)$
@@
$\hskip 28.45274pt-\dfrac{(n_{123}-s)(n_{123}-s-1)}{2}+\dfrac{(n_{12}-t)(n_{12}-t-1)}{2}+\dfrac{c_{3}(c_{3}-1)+c_{4}(c_{4}-1)}{2}$
$\check{\vartheta}^{*}_{t,s}=-s\dfrac{(2n_{123}-t-s-1)(2n_{12}-s-t)(s+2c_{4}-1)(2n_{123}-s)(2n_{1234}-s-1)}{4(2n_{123}-2s+1)(2n_{123}-2s-1)(n_{123}-s)^{2}}$
$C_{234}|t,s\rangle$ $=$ $\hat{\vartheta}_{t,s}|t,s+1\rangle+\hat{\psi}_{t,s}|t+1,s+1\rangle$
$+\vartheta_{t,s}|t,s\rangle+\psi_{t,s}|t+1,s\rangle$
$+\check{\vartheta}_{t,s}|t,s-1\rangle+\check{\psi}_{t,s}|t+1,s-1\rangle$
$\hat{\vartheta}_{t,s}=(s-t)(2n_{23}-s-t-1)$
@@
$\hskip 28.45274pt-\dfrac{(n_{123}-s)(n_{123}-s-1)}{2}+\dfrac{(n_{23}-t)(n_{23}-t-1)}{2}+\dfrac{c_{1}(c_{1}-1)+c_{4}(c_{4}-1)}{2}$
@@
$\hskip 19.91692pt=s\dfrac{(-2c_{1}+s-t)(2n_{123}-t-s-1)(s+2c_{4}-1)(2n_{123}-s)(2n_{1234}-s-1)}{4(2n_{123}-2s+1)(2n_{123}-2s-1)(n_{123}-s)^{2}}$
$\hat{\psi}_{t,s}=1$
$\psi_{t,s}=\dfrac{1}{2}+\dfrac{n_{123}(n_{123}+2c_{4}-1)}{2(n_{123}-s)(n_{123}-s-1)}$
$\check{\psi}_{t,s}=\dfrac{s(s+2c_{4}-1)(2n_{123}-s)(2n_{1234}-s-1)}{4(2n_{123}-2s+1)(2n_{123}-2s-1)(n_{123}-s)^{2}}$

### v2

@@ A.2 Case $\boldsymbol{\{P_{jk},D_{ijk},D_{jk\ell}\}}$
$[D_{ijk},[D_{jk\ell},P_{jk}]]$
$\qquad=[[P_{jk},D_{jk\ell}],D_{ijk}]=[(P_{k\ell}-P_{j\ell})P_{jk}+2D_{jk\ell}-2P_{j}P_{k\ell}+2P_{k}P_{j\ell},D_{ijk}]$
$\qquad=(P_{ik}P_{j\ell}+P_{ij}P_{k\ell}-2P_{i\ell}P_{jk})P_{jk}$
$\phantom{\qquad=}{}+(P_{k\ell}-P_{j\ell})\left((P_{ik}-P_{ij})P_{jk}+2D_{ijk}-2P_{j}P_{ik}+2P_{k}P_{ij}\right)$
$\phantom{\qquad=}{}-2(D_{ji\ell}+D_{ki\ell})P_{jk}-2P_{j}[P_{\ell k},D_{kij}]+2P_{k}[P_{\ell j},D_{jki}]$
$\qquad=-2P_{i\ell}P_{jk}^{2}+(P_{k\ell}P_{ik}+P_{j\ell}P_{ij})P_{jk}+2(P_{k\ell}-P_{j\ell})D_{ijk}-2D_{i\ell j}P_{jk}-2D_{i\ell k}P_{jk}$
$\phantom{\qquad=}{}-2P_{j}\left(P_{k\ell}P_{ij}-P_{i\ell}P_{jk}\right)+2P_{k}\left(P_{i\ell}P_{jk}-P_{j\ell}P_{ij}\right).$
@@
$+2P_{j}\left(P_{ik}P_{j\ell}-P_{i\ell}P_{jk}\right)-2P_{k}\left(P_{i\ell}P_{jk}-P_{ij}P_{j\ell}\right).$ (A.4)
Notice that the triple of $P$’s in (A.2) and (A.4) will sum together to give $2D_{\ell ki}P_{jk}+2D_{\ell ji}P_{jk}$ which will cancel with subsequent terms.
Putting everything together, we arrive at
@@ A.3 Case $\boldsymbol{\{P_{ij},D_{ijk},D_{jk\ell}\}}$
The first term on the right has not been involved in any of the previous computations as all of the previous computations have decreased the order of terms to at most triples of $P$’s, or equivalently a $P$ and $D$. We can remove this term by taking the commutator of a $PD$ relation (2.23) (again with the opposite ordering) with a $P$ as in
$[P_{jk},2D_{jk\ell}P_{i}+D_{k\ell i}P_{ij}+D_{\ell ji}P_{ik}+D_{jki}P_{i\ell}]=0,$
@@
$\qquad{}+2D_{k\ell i}D_{kji}+2D_{\ell ji}D_{jki}=0.$
Thus, after a permutation of indices, we obtain
@@
$[P_{jk},2D_{jk\ell}P_{i}+D_{k\ell i}P_{ij}+D_{\ell ji}P_{ik}+D_{jki}P_{i\ell}]={\rm In}+{\rm Out},$
with $\rm In$ and ${\rm Out}$ a collection of terms involving inner and outer quadratic relations, respectively, that is,
${\rm In}=-P_{ij}[P_{jk},D_{jk\ell}]+[P_{i\ell}[P_{jk},D_{jki}]]+P_{j\ell}[P_{ik},D_{ikj}]-[P_{ij},D_{ij\ell}]P_{jk},$
${\rm Out}=\left([P_{ji},D_{i\ell k}]+[P_{\ell i},D_{ijk}]-[P_{ij},D_{jk\ell}]\right)P_{jk}+\left([P_{jk},D_{ki\ell}]+[P_{ik},D_{jk\ell}]\right)P_{ij}$
$\phantom{Out=}{}+\left([P_{kj},D_{j\ell i}]+[P_{\ell j},D_{jik}]\right)P_{ik}+P_{ik}[P_{ij},D_{jk\ell}]-2P_{j}[P_{ik},D_{k\ell j}].$
Beginning with the inner relation terms, we can use the lower-order Jacobi identity proven earlier $[P_{i\ell}[P_{jk},D_{jki}]]=[P_{jk},[P_{i\ell},D_{jki}]]$, to simplify the term in ${\rm In}$,
$[P_{i\ell}[P_{jk},D_{jki}]]={}$ $[P_{jk},P_{ij}P_{k\ell}-P_{ik}P_{j\ell}]=P_{jk}P_{ij}P_{k\ell}-P_{jk}P_{j\ell}P_{ik}-(P_{ij}P_{k\ell}-P_{ik}P_{j\ell})P_{jk}$
[…]
The other 3 terms from ${\rm In}$ give
@@
$\qquad=P_{ij}(P_{j\ell}P_{jk}-P_{jk}P_{k\ell}-2P_{k}P_{j\ell}+2P_{j}P_{k\ell})+P_{j\ell}(P_{ik}P_{jk}-P_{ij}P_{ik}-2P_{i}P_{jk}+2P_{k}P_{ij})$
$\phantom{\qquad=}{}+(P_{i\ell}P_{ij}-P_{ij}P_{j\ell}-2P_{j}P_{i\ell}+2P_{i}P_{j\ell})P_{jk}$
$\qquad=(P_{ik}P_{j\ell}-P_{ij}P_{k\ell})P_{jk}-2P_{ij}D_{jk\ell}+(P_{i\ell}P_{jk}-P_{ik}P_{j\ell})P_{ij}+2P_{i\ell}D_{ijk}-2P_{j\ell}D_{jik}$
$\phantom{\qquad=}{}+4P_{k}D_{\ell ji}+2P_{j}(P_{ij}P_{k\ell}-P_{i\ell}P_{jk}).$
@@
Moving on to the ${\rm Out}$ term. We use Lemma 2.9 several times to simplify
${\rm Out}=[P_{jk},D_{k\ell i}]P_{jk}+[P_{ij},D_{j\ell k}]P_{ij}-[P_{ij},D_{jk\ell}]P_{ik}+P_{ik}[P_{ij},D_{jk\ell}]-2P_{j}[P_{ik},D_{k\ell j}].$
@@
${\rm Out}=[P_{jk},D_{k\ell i}]P_{jk}+[P_{ij},D_{j\ell k}]P_{ij}+2D_{ki\ell}P_{jk}+2P_{i\ell}D_{ikj}-2P_{j}[P_{ik},D_{k\ell j}].$
Hence, we have shown ${\rm In}+{\rm Out}=0$ and the Jacobi identity holds for these generators without additional constraints.
A.4 Case $\boldsymbol{\{D_{ij\ell},D_{ijk},D_{jk\ell}\}}$
The final case will be for 3 distinct $D$’s. Without loss of generality we chose the indices as above and consider the quantity
$[D_{ij\ell},[D_{ijk},D_{jk\ell}]]+[D_{ijk},[D_{jk\ell},D_{ij\ell}]]+[D_{jk\ell},[D_{ij\ell},D_{ijk}]].$
@@ A.4 Case $\boldsymbol{\{D_{ij\ell},D_{ijk},D_{jk\ell}\}}$
Appendix B Coefficients of the action of the proposed representation
@@
$n_{M}=N+\sum_{i\in M}c_{i},\qquad M\subseteq\{1,2,3,4\},$
$C_{1}|t,s\rangle=c_{1}(c_{1}-1)|t,s\rangle,\qquad C_{2}|t,s\rangle=c_{2}(c_{2}-1)|t,s\rangle,\qquad C_{3}|t,s\rangle=c_{3}(c_{3}-1)|t,s\rangle,$
$C_{4}|t,s\rangle=c_{4}(c_{4}-1)|t,s\rangle,\qquad C_{1234}|t,s\rangle=n_{1234}(n_{1234}-1)|t,s\rangle,$
$C_{12}|t,s\rangle=\varphi_{t,s}|t-1,s\rangle+\theta_{t,s}|t,s\rangle,$
$\varphi_{t,s}=(s-t)(N+1-t)(N+2c_{2}-t)(2n_{123}-t-s-1),$
$\theta_{t,s}=(n_{12}-t)(n_{12}-t-1),$
$C_{23}|t,s\rangle=\theta^{*}_{t,s}|t,s\rangle+|t+1,s\rangle,\qquad\theta^{*}_{t,s}=(n_{23}-t)(n_{23}-t-1),$
$C_{123}|t,s\rangle=\nu_{t,s}|t,s\rangle,\qquad\nu_{t,s}=(n_{123}-s)(n_{123}-s-1),$
$C_{34}|t,s\rangle=\hat{\phi}^{*}_{t,s}|t-1,s+1\rangle+\hat{\vartheta}^{*}_{t,s}|t,s+1\rangle+\phi^{*}_{t,s}|t-1,s\rangle+\vartheta^{*}_{t,s}|t,s\rangle+\check{\phi}^{*}_{t,s}|t-1,s-1\rangle$
${}+\check{\vartheta}^{*}_{t,s}|t,s-1\rangle,$
$\hat{\phi}^{*}_{t,s}=-(s-t)(s-t+1)(N+1-t)(N+2c_{2}-t),$
$\phi^{*}_{t,s}=\varphi_{t,s}(1-\psi_{t,s})=\left(\dfrac{1}{2}-\dfrac{n_{123}(n_{123}+2c_{4}-1)}{2(n_{123}-s)(n_{123}-s-1)}\right)$
×${}\times(s-t)(N+1-t)(N+2c_{2}-t)(2n_{123}-t-s-1),$
@@
${}=-s\dfrac{(s+2c_{4}-1)(2n_{123}-s)(2n_{1234}-s-1)(2n_{123}-t-s-1)}{4(2n_{123}-2s+1)(2n_{123}-2s-1)(n_{123}-s)^{2}}$
×${}\times(N+1-t)(2n_{123}-t-s)(N+2c_{2}-t),$
$\hat{\vartheta}^{*}_{t,s}=-(s-t)(s-2c_{3}-t+1),$
@@
${}-\dfrac{(n_{123}-s)(n_{123}-s-1)}{2}+\dfrac{(n_{12}-t)(n_{12}-t-1)}{2}+\dfrac{c_{3}(c_{3}-1)+c_{4}(c_{4}-1)}{2},$
$\check{\vartheta}^{*}_{t,s}=-s\dfrac{(2n_{123}-t-s-1)(2n_{12}-s-t)(s+2c_{4}-1)(2n_{123}-s)(2n_{1234}-s-1)}{4(2n_{123}-2s+1)(2n_{123}-2s-1)(n_{123}-s)^{2}},$
$C_{234}|t,s\rangle=\hat{\vartheta}_{t,s}|t,s+1\rangle+\hat{\psi}_{t,s}|t+1,s+1\rangle+\vartheta_{t,s}|t,s\rangle+\psi_{t,s}|t+1,s\rangle$
$\hat{\vartheta}_{t,s}=(s-t)(2n_{23}-s-t-1),$
@@
${}-\dfrac{(n_{123}-s)(n_{123}-s-1)}{2}+\dfrac{(n_{23}-t)(n_{23}-t-1)}{2}+\dfrac{c_{1}(c_{1}-1)+c_{4}(c_{4}-1)}{2},$
@@
${}=s\dfrac{(-2c_{1}+s-t)(2n_{123}-t-s-1)(s+2c_{4}-1)(2n_{123}-s)(2n_{1234}-s-1)}{4(2n_{123}-2s+1)(2n_{123}-2s-1)(n_{123}-s)^{2}},$
$\hat{\psi}_{t,s}=1,\qquad\psi_{t,s}=\dfrac{1}{2}+\dfrac{n_{123}(n_{123}+2c_{4}-1)}{2(n_{123}-s)(n_{123}-s-1)},$
$\check{\psi}_{t,s}=\dfrac{s(s+2c_{4}-1)(2n_{123}-s)(2n_{1234}-s-1)}{4(2n_{123}-2s+1)(2n_{123}-2s-1)(n_{123}-s)^{2}}.$
Acknowledgements
The authors would like to thank Nicolas Crampé and Éric Ragoucy for the fruitful discussion on the subject. The authors would also like to thank the anonymous referees for their comments and suggestions. SP would like to acknowledge the Simons Foundation Collaboration grant #3192112 as well as support as a CRM-Simons professorship and the hospitality of the CRM for a fruitful visit while much of the research was conducted.